. As the Manager of Walker Jones Memorial Hospital, you projected financial outcomes for 5 medical units in the hospital known for adding up to $ 30,000,000 all together of revenue to the hospital. The total cost of operating these units is $ 120,000 and a constant variable cost of $ 20,000. Consider that, at some point, each unit offered its regular patients lower medical costs of $ 30,000 off their bills. Calculate the Break-Even Number of Units.

4. From you answer above, prove that the operating income is 0 at break even.

5. As a manager, explain the implication of this outcome for the fiscal year.
Answer #1

Revenue to the hospital = 30,000,000

fixed cost = 120000

variable cost per unit = 20000

break even number of units = fixed cost/contribution margin per unit = 120000/((30000000-20000*5)/5)

Therefore break even number of units = 0.02

operating income = number of units * contribution margin - fixed cost = 0.02*(30000000-20000*5)/5 - 120000 =0

The implication from the break even number of units is, the hospital is achieving it's break even very easily even one unit is operating, it is achieving break even